OpenStudy (anonymous):

-1.5x + 2.5y + 3 = 0 6x - 10y + 6 = 0 Which statement about the pair of equations is true? A) The equation pair has no solutions. B) The equation pair has many solutions. C) The equation pair has exactly 1 solution. D) The equation pair has exactly 2 solutions.

OpenStudy (anonymous):

A) The equation has no solutions: because when u solve for y using second eqn and then substitute tht value of y in first eqn u dont get any result for x

OpenStudy (radar):

Multiplying the first equation by 4 gets -6x+10y+12=0 adding the second equation gets this: 6x -10y +6 =0 --------------- 0 + 0 +18=0 18=0 inconsistent equations with no solution for the system.